A rectangular garden has a perimeter of 500 feet, If one side of the garden has length x, then write a polynomial expression that gives its area. Multiply
this expression completely.

Answer by charu91(7)

2(l+b)=perimeter of rectangle = 500
length,l=x (given)
2(x+b)=500
x+b=250
b=250-x
area=l*b
=x*(250-x)
=250x-x^2
x^2-250x+area=0